Find the general solution of the following differential equations using method of undetermined coefficients:, (v) y''+2y'+2y =e^x cos2x.


1
Expert's answer
2022-03-02T11:09:52-0500

Corresponding homogeneous differential equation


"y''+2y'+2y =0"

Characteristic (auxiliary) equation


"r^2+2r+2=0"

"r_1=-1-i,r_2=-1+i"

The general solution of the homogeneous differential equation is


"y_h=c_1e^{-x}\cos x+c_2e^{-x}\sin x"

Find the particular solution of the non homogeneous differential equation


"y_p=Ae^x\cos 2x+Be^x\sin 2x"

"y_p'=Ae^x\cos 2x-2Ae^x\sin 2x"

"+Be^x\sin 2x+2Be^x\cos 2x"

"y_p''=Ae^x\cos 2x-2Ae^x\sin 2x"


"-2Ae^x\sin 2x-4Ae^x\cos 2x"

"+Be^x\sin 2x+2Be^x\cos 2x"

"+2Be^x\cos 2x-4Be^x\sin 2x"

Substitute


"-3Ae^x\cos 2x-4Ae^x\sin 2x-3Be^x\sin 2x"

"+4Be^x\cos 2x+2Ae^x\cos 2x-4Ae^x\sin 2x"

"+2Be^x\sin 2x+4Be^x\cos 2x+2Ae^x\cos 2x"

"+2Be^x\sin 2x=e^x\cos 2x"

"A+8B=1"

"-8A+B=0"

"A=1\/65, B=8\/65"


The general solution of the non homogeneous differential equation is


"y=c_1e^{-x}\cos x+c_2e^{-x}\sin x""+\dfrac{1}{65}e^x\cos 2x+\dfrac{8}{65}e^x\sin 2x"